a SHIP OF MASS 3*10^7 KG INITIALLY AT REST IT PULLED BY A FORCE 5*10^4n THROUGH A DISTANCE OF 3M aSSUME THAT THE RESISTANCE DUE TO WATER IN NEGLIGIBLE THE SPEED OF THE SHIP IS

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

acceleration of ship a=5*104 N/ 3*107 KG

                               =5/3 *10-3 m/sec2

speed of the ship after movieng 3 m 

v2 =u2 +2aS

    =0 +2* 5/3 *10-3   * (3)

    = 10-2

v=.1 m/sec

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Soumyadeep
5 years ago
M=3*10^7
F=5*10^4 N
S=3m
U=0
V=?
We know,
F=M.A
 
5*10^4=3*10^7*a
a=5/3*10^-3
We know,
 V²=U²+2as
V²=2*5/3*10-³ .3+0²
V²=10*10-³
V=√10²
V=.1m
            ANSWER..
 
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

acceleration of ship a=5*104 N/ 3*107 KG
=5/3 *10-3 m/sec2
speed of the ship after movieng 3 m
v2 =u2 +2aS
=0 +2* 5/3 *10-3 * (3)
= 10^-2
v=.1 m/sec

Thanks and Regards